OpenStudy (baileycali):

Write the equation of a line with a slope of 4 and a y-intercept of −3. 4x + y = −3 4x – y = −3 y = −3x + 4 y = 4x − 3

OpenStudy (mathmate):

The slope is given by the coefficient of x divided by the coefficient of y. If x and y are on the same side, switch the sign. For example, 2x+3y=7 slope = -(2/3) negative because x and y are on the same side x=-7y-5 slope = 1/(-7) = -1/7, (-7) because coefficient of y is -7. 4x + y = −3 : slope = -(4/1) = -4 (negative because x & y are on the same side The y-intercept can be found by solving the equation of y by putting x=0 Example: 5x+8y=4, put x=0, then 8y=4, y-intercept=4/8=1/2 4x + y = −3, put x=0, then y=-3 = y-intercept Proceed the same way with the other choices to find your answer.